# Model Editing Harms General Abilities of Large Language Models: Regularization to the Rescue

### v3

@@ 4.3 Results
Since single-editing can be regarded as a special case of sequential-editing when the number of edits is 1, this subsection mainly discussed instance- and sequential-editing. KN and ROME that support instance-editing but not batch-editing were adopted to facilitate this exploration. MEND and MEMIT that support batch- and sequential-editing will be explored later in this subsection. Figure 3 presents the performance on general tasks of edited models using KN or ROME to edit GPT-2 XL and LLaMA-1 (7B) as the number of edits increases. Due to limited space, readers can refer to Appendix C for the results of editing LLaMA-2 (7B) which show similar trends. It can be seen that although there is only one instance per editing operation, the performance of edited models on various tasks fluctuates significantly and shows a downward trend as the number of edits increases. Strikingly, the use of KN resulted in a drastic performance degradation to nearly zero on all selected tasks with just a single edit. These findings underscore two key insights. First, the selected LLMs are not robust to weight perturbations even if less than 1% of the parameters are edited, whereby slight perturbations may significantly affect their general abilities. Second, these outcomes also shed light on the challenging nature of effectively coupling current editing algorithms with LLMs. The difficulty lies in the dual objective of improving model factuality while simultaneously maintaining their general abilities. The observed trends indicate that existing editing algorithms face grand challenges in achieving this delicate balance, emphasizing the need for further research and development in the refinement of editing methodologies for LLMs.
@@
This subsection delved into batch- and single-editing to explore the impact of batch size for scaling up the editing scope. Only MEND and MEMIT that supported batch-editing were adopted to facilitate this exploration. Figure 4 presents the performance on general tasks of edited models using MEND or MEMIT to edit GPT-2 XL and LLaMA-1 (7B) with different batch sizes. Readers can refer to Appendix C for the results of editing LLaMA-2 (7B). Remarkably, even with only one single editing operation, edited models exhibited a trend of performance degradation as the batch size increases in most cases. This consistent decrease in performance underlines the sensitivity of the models to increases in batch size, emphasizing the significance of carefully scaling knowledge editing for optimal updates. Therefore, we call for more research work on scalable editing to facilitate efficient editing of multiple instances.
@@
In order to holistically take into account the interplay between batch size and sequential-editing, a joint setting of batch- and sequential-editing was explored to understand how these two factors collaboratively influence the overall performance of edited models. Figure 11 to Figure 16 in Appendix C present the performance of using MEND or MEMIT to edit GPT-2 XL, LLaMA-1 (7B) and LLaMA-2 (7B) respectively as the number of edits increases. These results also echo our observations on sequential-editing, and those on batch-editing respectively.
@@ 4.4 Analysis of Causes of Side Effects
We first show how the weights change in instance- and single-editing. The statistics show that only 20% of the elements in the update weight $\Delta W$ have $\delta$ greater than 0.077, while only 10% of the elements have $\delta$ greater than 0.171. These results are averaged by 100 random single edits in the ZsRE dataset. It can be seen that the update weight $\Delta W$ might be quite sparse, while most elements in $\Delta W$ are minor. Manhattan distance between the updated and original weights is also calculated as a measurement of distinction.
@@ 5 RECT: RElative Change in weighT
Figure 6: Comparison of (a) non-regularization and (b) the proposed RECT regularization. Elements in green denote the top-kthat change the most according to $\delta$, and are considered as the principle editing information. k = 25 in this figure for illustration.
@@ 5.1 Approach
where function $f$ denotes the calculation method of update weight $\Delta W$ for different editing methods, e.g., ROME (Meng et al., 2022).
@@ 5.3 Results of RECT
(a) Summarization
(b) Open-domain QA
(c) Closed-domain QA
(d) Sentiment Analysis
The effectiveness of a regularization method should be illustrated from two perspectives. First, regularizing the edit update weight should not harm its editing performance, i.e., edited models should still remember the new editing facts and generalize to related facts. Second, the regularized edited models should be able to preserve the general abilities compared with unregularized ones. We show the results of regularizing ROME here and readers can refer to Appendix C.2 for more results of regularizing other editing methods on more LLMs and downstream tasks.
(a) GPT-2 XL
(b) LLaMA-1 (7B)
@@
Figure 8 presents the comparison of introducing various regularization methods to edit GPT-2 XL and LLaMA-1 (7B). From these results we can have the following findings. First, compared with unregularized $\Delta W$, RECT that keeps the original values of an appropriate amount of top-40% elements in $\Delta W$ and sets the remaining elements to zero can help maintain over 94% majority of reliability and generalization, and even improve locality. It is natural that reliability and generalization slightly drop when setting partial elements to zero since partial editing information is removed. The reason why locality is improved is probably because those elements with low $\delta$ corresponding to some noise and outliers in the editing data are removed to prevent from editing overfitting, so the edited models are more robust. However, setting excessive elements in $\Delta W$ to zero, e.g., RECT top-20%, might hurt the editing performance as partial important editing information is accidentally removed. Furthermore, compared with Random 40% and PCA 40%, RECT top-40% achieves the best performance, indicating its effectiveness in selecting the most principal editing information. It is notable that RECT also exhibits advantages in terms of efficiency, since it eliminates the complex calculations required in PCA.
@@
Figure 7 presents the comparison of introducing various regularization methods to edit GPT-2 XL. As the proportion of elements in $\Delta W$ set to 0 increases, the more editing overfitting is regularized, the smaller the change to the original weight, so the general abilities can be more preserved. Results show that regularized edited models are able to preserve the general abilities compared with unregularized ones in most tasks such as summarization, open- and closed-domain QA. It is worth noting that it still poses a challenge for some tasks such as sentiment analysis, and remains unclear whether it works for larger number of edits, which will be left to future work.
@@ 6 Conclusion
Model updating technology has been catalyzing the continuous iteration of advanced and trustworthy LLMs. This paper studies model editing and for the first time raises concerns whether model editing has any side effects. The systematical evaluation reveals that current methods unintentionally hurt the general abilities of LLMs no matter in instance- or batch-editing, and single- or sequential-editing. Our analysis of the causes reveals that model editing results in overfitting to the editing facts, and the accumulation of overfitting across multiple edits can amplify the negative impact. The proposed RECT regularization method has been proven to effectively prevent overfitting of new editing facts, thus preserving both the editing and general downstream task performance.
@@
Appendix A Task Prompts
@@
Appendix B Editing Methods & Datasets
@@
Four popular editing methods as compared in Table 3 were selected to measure their performance in improving factuality as well as their impairment in the general abilities of LLMs, including:
KN (Dai et al., 2022) [^1] https://github.com/EleutherAI/knowledge-neurons first selected neurons that were associated with knowledge expression via gradient-based attributions, and then modified MLP layer at the rows corresponding to those neurons by adding scaled embedding vectors.
MEND (Mitchell et al., 2022a) [^2] https://github.com/eric-mitchell/mend learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form.
ROME (Meng et al., 2022) [^3] https://github.com/kmeng01/rome first localized the factual knowledge at a specific layer in the transformer MLP modules, and then updated the knowledge by directly writing new key-value pairs in the MLP module.
MEMIT (Meng et al., 2023) [^4] https://github.com/kmeng01/memit expanded the capabilities of ROME by enabling the editing of large amounts of factual data through the updating of a sequence of MLP layers.
All experiments were conducted using the EasyEdit tool (Wang et al., 2023), ensuring standardized and reproducible evaluation. All editing instances were randomly sampled from the editing dataset.
The popular model editing dataset Zero-Shot Relation Extraction (ZsRE) (Levy et al., 2017) used in previous work (Cao et al., 2021; Meng et al., 2022; Yao et al., 2023) was adopted in our experiments. ZsRE is a QA dataset using question rephrasings generated by back-translation as the equivalence neighborhood. Each input is a question about an entity, and plausible alternative edit labels are sampled from the top-ranked predictions of a BART-base model trained on ZsRE.
Appendix C Extensive Evaluation Results
C.1 Results of Side Effects of Model Editing
@@ C.2 Results of RECT
(b) MEMIT on GPT-2 XL
(c) MEMIT on LLaMA-1 (7B)
(d) MEMIT on LLaMA-2 (7B)

### v4

@@ 4.3 Results
Since single-editing can be regarded as a special case of sequential-editing when the number of edits is 1, this subsection mainly discussed instance- and sequential-editing. KN and ROME that support instance-editing but not batch-editing were adopted to facilitate this exploration. MEND and MEMIT that support batch- and sequential-editing will be explored later in this subsection. Figure 3 presents the performance on general tasks of edited models using KN or ROME to edit GPT-2 XL and LLaMA-1 (7B) as the number of edits increases. Due to limited space, readers can refer to Appendix D for the results of editing LLaMA-2 (7B) which show similar trends. It can be seen that although there is only one instance per editing operation, the performance of edited models on various tasks fluctuates significantly and shows a downward trend as the number of edits increases. Strikingly, the use of KN resulted in a drastic performance degradation to nearly zero on all selected tasks with just a single edit. These findings underscore two key insights. First, the selected LLMs are not robust to weight perturbations even if less than 1% of the parameters are edited, whereby slight perturbations may significantly affect their general abilities. Second, these outcomes also shed light on the challenging nature of effectively coupling current editing algorithms with LLMs. The difficulty lies in the dual objective of improving model factuality while simultaneously maintaining their general abilities. The observed trends indicate that existing editing algorithms face grand challenges in achieving this delicate balance, emphasizing the need for further research and development in the refinement of editing methodologies for LLMs.
@@
This subsection delved into batch- and single-editing to explore the impact of batch size for scaling up the editing scope. Only MEND and MEMIT that supported batch-editing were adopted to facilitate this exploration. Figure 4 presents the performance on general tasks of edited models using MEND or MEMIT to edit GPT-2 XL and LLaMA-1 (7B) with different batch sizes. Readers can refer to Appendix D for the results of editing LLaMA-2 (7B). Remarkably, even with only one single editing operation, edited models exhibited a trend of performance degradation as the batch size increases in most cases. This consistent decrease in performance underlines the sensitivity of the models to increases in batch size, emphasizing the significance of carefully scaling knowledge editing for optimal updates. Therefore, we call for more research work on scalable editing to facilitate efficient editing of multiple instances.
@@
In order to holistically take into account the interplay between batch size and sequential-editing, a joint setting of batch- and sequential-editing was explored to understand how these two factors collaboratively influence the overall performance of edited models. Figure 11 to Figure 16 in Appendix D present the performance of using MEND or MEMIT to edit GPT-2 XL, LLaMA-1 (7B) and LLaMA-2 (7B) respectively as the number of edits increases. These results also echo our observations on sequential-editing, and those on batch-editing respectively.
4.4 Analysis of Causes of Side Effects
@@
We first show how the weights change in instance- and single-editing. Typically, one editing operation is to add an edit update weight $\Delta W$ to the original weight $W$, where $\Delta W$ is calculated aiming to insert new editing facts. Here, we define the absolute value of the relative change in weight $\delta=|\frac{\Delta W}{W}|$ to characterize the degree of change of each element in the update weight $\Delta W$. The statistics show that only 20% of the elements in the update weight $\Delta W$ have $\delta$ greater than 0.077, while only 10% of the elements have $\delta$ greater than 0.171. These results are averaged by 100 random single edits in the ZsRE dataset. It can be seen that the update weight $\Delta W$ might be quite sparse, while most elements in $\Delta W$ are minor. Manhattan distance between the updated and original weights is also calculated as a measurement of distinction.
@@ 5.1 Approach
Figure 6: Comparison of (a) non-regularization and (b) the proposed RECT regularization. Elements in green denote the top-kthat change the most according to $\delta$ and are considered as the principle editing information,which should keep their original values. k = 25 in this figure for illustration.
(a) ROME on GPT-2 XL
(b) ROME on LLaMA-1 (7B)
(c) MEMIT on GPT-2 XL
(d) MEMIT on LLaMA-1 (7B)
(a) Summarization
(b) Open-domain QA
(c) Closed-domain QA
(d) Sentiment Analysis
@@
where function $f$ denotes the calculation method of update weight $\Delta W$ for different editing methods, e.g., ROME Meng et al. (2022).
@@ 5.3 Results of RECT
The effectiveness of a regularization method should be illustrated from two perspectives. First, regularizing the edit update weight should not harm its editing performance, i.e., edited models should still remember the new editing facts and generalize to related facts. Second, the regularized edited models should be able to preserve the general abilities compared with unregularized ones.
@@
Figure 7 presents the the results of regularizing ROME or MEMIT on GPT-2 XL or LLaMA-1 (7B). Readers can refer to Appendix D.2 for the results on LLaMA-2 (7B). From these results we can have the following findings. First, compared with unregularized $\Delta W$, RECT that keeps the original values of an appropriate amount of top-40% elements in $\Delta W$ and sets the remaining elements to zero can help maintain over 94% majority of reliability and generalization, and even improve locality. It is natural that reliability and generalization slightly drop when setting partial elements to zero since partial editing information is removed. The reason why locality is improved is probably because those elements with low $\delta$ corresponding to some noise and outliers in the editing data are removed to prevent from editing overfitting, so the edited models are more robust. However, setting excessive elements in $\Delta W$ to zero, e.g., RECT top-20%, might hurt the editing performance as partial important editing information is accidentally removed. Furthermore, compared with Random 40% and PCA 40%, RECT top-40% achieves the best performance, indicating its effectiveness in selecting the most principal editing information. It is notable that RECT also exhibits advantages in terms of efficiency, since it eliminates the complex calculations required in PCA.
@@
Figure 8 presents how the downstream task performance change with respect to introducing various regularization methods to edit GPT-2 XL. Readers can refer to Appendix D.2 for the results on more downstream tasks. From these results we can have the following findings. As the proportion of elements in $\Delta W$ set to 0 increases, the more editing overfitting is regularized, the smaller the change to the original weight, so the general abilities can be more preserved. Results show that regularized edited models are able to preserve the general abilities compared with unregularized ones in most tasks such as summarization, open- and closed-domain QA. It is worth noting that it still poses a challenge for some tasks such as sentiment analysis, and remains unclear whether it works for larger number of edits, which will be left to future work.
@@ 6 Conclusion
Model updating technology has been catalyzing the continuous iteration of advanced and trustworthy LLMs. This paper studies model editing and for the first time raises concerns whether model editing has any side effects on the general abilities of LLMs. The systematical evaluation reveals that current methods unintentionally hurt the general abilities of LLMs no matter in instance- or batch-editing, and single- or sequential-editing. Our analysis of the causes reveals that model editing results in overfitting to the editing facts, and the accumulation of overfitting across multiple edits can amplify the negative impact. The proposed RECT regularization method has been proven to effectively prevent overfitting of new editing facts, thus preserving both the editing and general downstream task performance.
@@
Acknowledgement
This research is based upon work supported by an Amazon AGI foundation research award, a google research scholar grant, CISCO sponsored research award, and NSF #2331966. We thank Tanmay Parekh, Po-Nien Kung, Sidi Lu, Fabrice Harel-Canada, UCLA NLP group members and anonymous reviewers for their valuable feedback.
Appendix A Details of Evaluation Metrics
Reliability
Generalization
Locality
Appendix B Task Prompts
@@
Appendix C Editing Methods & Datasets
@@
Four popular editing methods as compared in Table 3 were selected to measure their performance in improving factuality as well as their impairment in the general abilities of LLMs, including: KN Dai et al. (2022) [^2] https://github.com/EleutherAI/knowledge-neurons first selected neurons that were associated with knowledge expression via gradient-based attributions, and then modified MLP layer at the rows corresponding to those neurons by adding scaled embedding vectors. MEND Mitchell et al. (2022a) [^3] https://github.com/eric-mitchell/mend learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form. ROME Meng et al. (2022) [^4] https://github.com/kmeng01/rome first localized the factual knowledge at a specific layer in the transformer MLP modules, and then updated the knowledge by directly writing new key-value pairs in the MLP module. MEMIT Meng et al. (2023) [^5] https://github.com/kmeng01/memit expanded the capabilities of ROME by enabling the editing of large amounts of factual data through the updating of a sequence of MLP layers.
All experiments were conducted using the EasyEdit tool Wang et al. (2024a), ensuring standardized and reproducible evaluation. All editing instances were randomly sampled from the editing dataset.
The popular model editing dataset Zero-Shot Relation Extraction (ZsRE) Levy et al. (2017) used in previous work Cao et al. (2021); Meng et al. (2022); Yao et al. (2023) was adopted in our experiments. ZsRE is a QA dataset using question rephrasings generated by back-translation as the equivalence neighborhood. Each input is a question about an entity, and plausible alternative edit labels are sampled from the top-ranked predictions of a BART-base model trained on ZsRE.
Appendix D Extensive Evaluation Results
D.1 Results of Side Effects of Model Editing
@@ D.2 Results of RECT
(b) MEMIT on LLaMA-2 (7B)